OpenStudy (anonymous):

How do you write cos(6x) in terms of cosx and sinx?

OpenStudy (lgbasallote):

im seeing this as \[\cos [2(3x)]\]

OpenStudy (lgbasallote):

let a = 3x to make it easier \[\cos 2a\] do you see it?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

@lgbasallote Please continue...

OpenStudy (experimentx):

\[ (\cos 6x + i \sin 6x) = (\cos x + i \sin x)^6 \\ \text{expand it binomially and get it's real part.}\]

OpenStudy (anonymous):

still need help? @shivraj

OpenStudy (anonymous):

no thanks